OpenStudy (anonymous):

don't really understand this. The sum of two consecutive even integers is 226. What is the larger integer?

OpenStudy (anonymous):

do you know what "consecutive" means?

OpenStudy (anonymous):

x+(x+1)=226

OpenStudy (anonymous):

no acually.

OpenStudy (anonymous):

for example 55 and 56 are consecutive guess and check

OpenStudy (anonymous):

i guess 111 and 112, but they add up only to 123, so too low try a larger pair

OpenStudy (anonymous):

that is what i thought, because if you realize "consecutive" means one integer then the next integer, you can do this more or less in your head

OpenStudy (anonymous):

oops it says "consecutive EVEN integers"

OpenStudy (anonymous):

like say 110 and 112

OpenStudy (anonymous):

but 110 ad 112 add up to 222, so try a larger pair

OpenStudy (anonymous):

114 and 115?

OpenStudy (anonymous):

(2n-2) + 2n = 226 or 4n -2 = 226 or 2n = 114 ← the larger integer